AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 194     Entry time: Mon Feb 14 09:56:39 2022
Author: CB 
Subject: Shifters instructions 2022 

EMERGENCY SHUTDOWNS

1. CARME motors: Q key on keyboard while on workspace 1 of appc218 (right of this double-monitor setup)

2. Internal target: Q on keyboard while on workspace 3 of appc218, go to server room, press "CLOSE" button on DUMP.

 

CARME DAQ

CARME computer has 6 workspaces (WS), two monitors for each WS.

  • WS1 : left monitor - fee power

          right monitor - fee power server

 

  • WS2 : left monitor - DAQ controls

          right monitor - CAEN bias control

Most important tabs (right hand monitor)

  • Run control : shows daq status, stopped / going
  • Fee temperatures : temperatures of fees labelled aida01, aida02, aida03, aida04, aida05
  • Statistics : shows the statistics for each fee
  • System wide checks : checks for the setup
  • Spectrum browser : loads histograms for fees and fee channels from preset layouts
  • Asic control : sets asic parameters

 

  • WS3 : left monitor - Merge control, tape server

          right monitor - servers, data being written

  • WS4 : leakage current plot
  • WS5 : online data monitoring
  • WS6 : browser, elog
  • Motors are running on second computer

 

TO DO FOR THE SHIFTERS - CHECK EVERY HOUR AND POST TO ELOG WITH SCREENSHOTS

 

-->IMPORTANT NOTE: MIDAS DAQ can respond slowly. DO NOT try to change tab or give new command while the previous one is being executed. Check activity in right-hand part of tab currently in use. Do not double click on buttons.

 

Motors computer (Computer to the right)

  • Check motors cycle is running. The code will cycle through the loaded motors instructions. When an instruction is being done it is highlighted in bold. Check the bold highlighting cycles through entire set of instructions.
  • Check potentiometer is reproducing same max/min values for each cycle of instructions. 74.02 fully in, 29.5 fully out.
  • Detectors are on LEFT arm. Nothing on RIGHT arm which should not move.
  • No need for screenshots, but record in the ELOG you checked this.

 

Merger, Tape server (WS3)

  • Check NewMerger tab. Hit reload button and see if 'current links with data' and 'current merged data rate' are updating. All links should go green, but not necessarily at the same time. Keep hitting refresh until you see all of them going green at least once.
  • Data rate should be around 100,000. Take screenshot.
  • Check tapeService tab. Hit reload button. Check data rate in Kbytes/sec box. Record data rate to elog. This is where the current run number is displayed. Take screenshot.
  • On the right-hand monitor, bottom right terminal, check the current run fragment (sub-run) is being written and is slowly increasing in size. It refreshed automatically every 5 seconds.

 

Leakage current (WS4)

  • Check leakage current plot is stable. Make sure it is updating and time is correct. Take screenshot. Check present values on WS2, CAEN window, channel 1.

 

DAQ (WS2)

  • Check Fee temperatures tab. Hit reload, temperatures will update and should all be green. 1-2 C above safe limit (top) is OK. More than that - warn expert shifter immediately. Take screenshot.
  • Check system wide checks tab. Click on 'Check Clock Status', 'Check ADC Calibration' and 'Check the White rabbit decoder status' to run checks. Copy-paste test in the ELOG the first time you do it. If there are no changes throughout the shift, just record no changes.
  • Check statistics tab. Hit Update All to see if the counter and rate are updating. Change counter being displayed using drop-down menu on the right hand side of the left monitor.
    • Good events - aida 1-4 should have similar counting rates. If any rate is a lot higher than the others, warn expert shifter.
    • ADC coincidence-info #8
    • Take screenshot of each for the ELOG
  • Spectrum browser tab. Check it is displaying aida 1-4 rate (top right of each histogram). IF NOT: select layout 1 from 'layout ID' then restore layout from 'Arrange functions' drop down menu. Select log scale and change xmax as 128. This displays the data rate per strip. If any strips are a lot higher than the others, or very different from previous entry, make a note in the ELOG. Take screenshot.

 

Online monitoring (WS 5)

  • Check the following plots are updating
    • LowEnergyXYTotal (right click - col - colz). This shows the total number of events, per pixel, since the code was started / re-started.
    • LowEnergyXYRate (right click - col - colz). This shows the rate over the last minute.
    • LowEnergyExTotal (double left click. Right click on y axis - Log scale on y). This shows the total energy deposited on all X strips since code restart.
    • LowEnergyEyTotal (as above). as above, Y strips.
  • Please take screenshot of each

 

Attachment 1: good_events_2022-02-15_09-40-40.png  103 kB  Uploaded Tue Feb 15 08:49:21 2022  | Hide | Hide all
good_events_2022-02-15_09-40-40.png
Attachment 2: ADC_data_items_2022-02-15_09-42-55.png  102 kB  Uploaded Tue Feb 15 08:49:37 2022  | Hide | Hide all
ADC_data_items_2022-02-15_09-42-55.png
Attachment 3: merger_control_2022-02-15_09-36-06.png  127 kB  Uploaded Tue Feb 15 08:49:43 2022  | Hide | Hide all
merger_control_2022-02-15_09-36-06.png
Attachment 4: leakage_current_2022-02-15_09-37-22.png  21 kB  Uploaded Tue Feb 15 08:49:49 2022  | Hide | Hide all
leakage_current_2022-02-15_09-37-22.png
Attachment 5: fee_temperatures_2022-02-15_09-38-36.png  111 kB  Uploaded Tue Feb 15 08:49:56 2022  | Hide | Hide all
fee_temperatures_2022-02-15_09-38-36.png
Attachment 6: spectrum_browser_2022-02-15_09-44-10.png  152 kB  Uploaded Tue Feb 15 08:50:06 2022  | Hide | Hide all
spectrum_browser_2022-02-15_09-44-10.png
Attachment 7: correlation_info8_2022-02-15_09-41-16.png  99 kB  Uploaded Tue Feb 15 08:50:13 2022  | Hide | Hide all
correlation_info8_2022-02-15_09-41-16.png
Attachment 8: lowEnergyExTotalDSSD0_2022-02-15_09-48-13.png  189 kB  Uploaded Tue Feb 15 08:50:19 2022  | Hide | Hide all
lowEnergyExTotalDSSD0_2022-02-15_09-48-13.png
Attachment 9: lowEnergyEyTotalDSSD0_2022-02-15_09-47-28.png  187 kB  Uploaded Tue Feb 15 08:50:26 2022  | Hide | Hide all
lowEnergyEyTotalDSSD0_2022-02-15_09-47-28.png
Attachment 10: lowEnergyXYTotalDSSD0_2022-02-15_09-45-35.png  226 kB  Uploaded Tue Feb 15 08:50:33 2022  | Hide | Hide all
lowEnergyXYTotalDSSD0_2022-02-15_09-45-35.png
ELOG V3.1.4-unknown